Let m,n be non negative integers and.let i, j element of N be even . let f:R2 to R be defined by f(0,0)=0 and

f(x,y) = (xmyn)/(xi+yj) for (x,y) not equal to (0,0) .show that f is continuous at (0,0) if and only if ,if mj+mi >ij

1
Expert's answer
2021-03-30T07:34:39-0400

Consider the path "x(t)=t^j, y(t)=kt^i", which approaches to (0,0) as t tends to 0.

If f(x,t) is a continuous at (0,0) then "f(x(t),y(t))=(1+k^j)^{-1}t^{mj+ni-ij}" must have a bounded limit as t tends to 0. Hence it must be "mj+ni-ij\geq 0" . and then

"\lim\limits_{t\to 0}f(x(t),y(t))=\begin{Bmatrix}
 0 & if\, mj+ni-ij> 0\\
 (1+k^j)^{-1} & if\, mj+ni-ij= 0
\end{Bmatrix}"

So, if "mj+ni-ij= 0" then the limit value varies depending on k, hence, "\lim\limits_{(x,y)\to (0,0)}f(x,y)" doesn't exist and the function may not be continuous at (0,0).

So, the necessity of condition "mj+ni-ij> 0" or, equivalently, "\frac{m}{i}+\frac{n}{j}>1" is proved.

Let us show now that this condition is sufficient for the function f(x,y) to be continuous at (0,0).

Let "s=\frac{m}{i}+\frac{n}{j}", "p=\frac{is}{m}", "q=\frac{js}{n}." Then "\frac{1}{p}+\frac{1}{q}=1".

Since i and j are even positive integers, "x^i+y^j=|x|^i+|y|^j" for every x,y.

By Young's inequality for products we have

"|x|^i+|y|^j=(|x|^{m/s})^p+(|y|^{n/s})^q=\frac{(p^{1/p}|x|^{m/s})^p}{p}+\frac{(q^{1/q}|y|^{n/s})^q}{q}\geq p^{1/p}q^{1/q}|x|^{m/s}|y|^{n/s}=C(|x|^m |y|^n)^{1/s}"

where "C=p^{1\/p}q^{1\/q}". So,

"|f(x,y)|=\frac{|x|^m|y|^n}{|x|^i+|y|^j}\leq C^{-1}(|x|^m|y|^n)^{1-1/s}\to 0"

Therefore, if "mj+ni-ij> 0" then "f(x,y)\to0=f(0,0)" as (x,y) tends to (0,0), hence, f(x,y) is continuous at (0,0).